OpenStudy (rossmeador):

Find the inverse of each function. y=log(base3)(x+8) y=-9log(basex)4

OpenStudy (anonymous):

x=y3-8

OpenStudy (rossmeador):

And elaborate, please.

OpenStudy (ujjwal):

y=log(base3)(x+8) therefore, 3^y=x+8 interchanging x and y 3^x=y+8 or, Y=3^x -8 so, that's your inverse function for 1st one.

OpenStudy (ujjwal):

Can you try the other one in the similar way?

OpenStudy (anonymous):

second is pre much the same 1/y = -1/9 log4 (x)

OpenStudy (anonymous):

suno...power raise karne ka ulta logarithm karna hota hai. samjhe?

OpenStudy (rossmeador):

What would I do with the -9 for the second one?

OpenStudy (ujjwal):

take it to the power. xlogy =logy^x

OpenStudy (rossmeador):

So it would be log(basex)4^-9?

OpenStudy (ujjwal):

yep

OpenStudy (rossmeador):

y=log(basex)4^-9 so x^y=4^-9?

OpenStudy (ujjwal):

yep!

OpenStudy (rossmeador):

And you interchange x and y, so y^x=4^-9 and solve for y and that's the answer?

OpenStudy (ujjwal):

yep!